OpenStudy (anonymous):

\[A=(2yz)i-(x ^{2}y)j+(xz ^{2})k,B=(x ^{2})i+(yz)j-(xy)k\]what is\[(B.\nabla)A\]

OpenStudy (turingtest):

\[(\vec B\cdot\nabla)\vec A\]?

OpenStudy (turingtest):

...as in the scalar multiple of \(\text{div}\vec B\) with \(\vec A\) ???

OpenStudy (anonymous):

should be, lol

OpenStudy (anonymous):

i cant obtain the answer provided by my tutor

OpenStudy (turingtest):

do you know how to get the divergence of \(\vec B\)\[\text{div}\vec B=?\]

OpenStudy (anonymous):

yea. i get \[(2x+z+0)\]

OpenStudy (anonymous):

or i get it wrongly LOL

OpenStudy (turingtest):

nope that's right

OpenStudy (anonymous):

then the scalar multiplication, i get\[(2yz(2x+z), -x ^{2}y(2x+z), 0)\]but the answer provided is \[(2y(z ^{2}-xy), -x ^{2y}(2x+z), x ^{2}z(z-2y))\]

OpenStudy (turingtest):

well that's not \[(\nabla\cdot\vec B)\vec A\]

OpenStudy (turingtest):

let's see if it'swell that's not \[(\nabla\cdot\vec A)\vec B\]real quick maybe?

OpenStudy (anonymous):

u mean the answer's has problem?

OpenStudy (turingtest):

yeah, if the question is\[(\text{div}\vec B)\vec A\]then the book made a mistake or something ...unless I've lost my mind, which happens from time to time at 3:30am

OpenStudy (anonymous):

if u dont mind, 1more question?

OpenStudy (turingtest):

sure, try me

OpenStudy (anonymous):

given\[\phi =2x ^{2}yz ^{3}\] then what is \[(A \times \nabla)\phi\]

OpenStudy (anonymous):

the cross product i get\[(z ^{2}-2y)j+(2z+2xy)k\], correct?

OpenStudy (turingtest):

is that just\[\vec A\times\nabla\]up there? 'cause that's right as far as I can see for that, but I don't see \(\phi\)

OpenStudy (anonymous):

yes its vector of A crossed with nabla, and outside the phi is given as\[2x ^{2}yz ^{3}\]

OpenStudy (anonymous):

T_T the test is 2mr and his tutorial aint helping anytg lol

OpenStudy (turingtest):

lol just multiply be \(\phi\) as a scalar now and I guess that's it kinda dumb problem if you ask me

OpenStudy (turingtest):

it want's the negative curl and a simple scalar multiplication just an exercise I guess

OpenStudy (anonymous):

i get \[2x ^{2}yz ^{3}(z ^{2}-2y)j+2x ^{2}yz ^{3}(2z+2xy)k\] the answer given is \[-2x ^{3}z ^{2}(3xy ^{2}+z ^{3})i+4x ^{2}yz ^{2}(z ^{2}-3y)j+4x ^{2}yz ^{3}(z+xy)k\]._.

OpenStudy (anonymous):

the i popped out from nowhere

OpenStudy (turingtest):

we are not understanding what they want it seems

OpenStudy (turingtest):

now the problem is to figure out what they are asking lol because it ain't\[(\vec A\times\nabla)\phi\]with phi as a scalar multiple

OpenStudy (turingtest):

like you said, the \(\hat i\) component is clearly zero after \(\nabla\times A\)

OpenStudy (anonymous):

i think i'll just leave those aside.. lol. mayb my tutor is blur afterall

OpenStudy (turingtest):

I feel bad I can't help, but something is amiss here better luck next time, promise as for me, it's goodnight :)

OpenStudy (anonymous):

kk nitez